A solid square rod is cantilevered at one end. The rod is 0.6 m long and supports a completely reversing transverse load at the other end of ±2 kN. The material is AISI 1080 hot-rolled steel. If the rod must support this load for 104 cycles with a factor of safety of 1.5 at 100°C, what dimension should the square cross section have? Neglect any stress concentrations at the support end.




1
Expert's answer
2022-03-22T23:36:03-0400

"L=0.6 m,F \na\n\u200b\n =2 kN,n=1.5,N=10 \n4\n cycles, S \nut\n\u200b\n =770 MPa,S \ny\n\u200b\n =420 MPa"

First evaluate the fatigue strength.

"S \ne\n\u2032\n\u200b\n =0.5(770)=385 MPa"

"S \ne\n\u2032\n\u200b\n =0.5(770)=385 MPa"

ince the size is not yet known, assume a typical value of "k_{b}\u200b = 0.85" and check later. All other modifiers are equal to one.


"S \ne\n\u200b\n ="

"=0.488(0.85)(385)=160 MPa"

"S \nut\n\u200b\n =770\/6.89=112 kpsi"


"f = 0.83"

"a=\frac{\left(f S_{u t}\right)^{2}}{S_{e}}=\frac{[0.83(770)]^{2}}{160}=2553 MPa"

"b=-\frac{1}{3} \log \left(\frac{f S_{u t}}{S_{e}}\right)=-\frac{1}{3} \log \left(\frac{0.83(770)}{160}\right)=-0.2005"

"S_{f}=a N^{b}=2553\left(10^{4}\right)^{-0.2005}=403 MPa"

Now evaluate the stress.

"M \nmax\n\u200b\n =(2000 N)(0.6 m)=1200 N\u22c5m"


"σ_a =σ _{max}=\frac{Mc}{I} = \frac{6M}{B^3}= \frac{7200}{b^3} Pa"


"\sigma_{max}= \frac{7200}{0.030^3}\times 10^-6= 267 Mpa"


"n_y= \frac{S_y}{\sigma_{max}}=\frac{420}{267}=1.57"